OpenStudy (anonymous):

Which set of polar coordinates describes the same location as the rectangular coordinates (-3,3)

OpenStudy (anonymous):

"which set..." implies there is a choice....

jimthompson5910 (jim_thompson5910):

Hints: r = sqrt(x^2+y^2) theta = arctan(y/x)

OpenStudy (anonymous):

A) (3,45degrees) B) (3 squareroot 2, 315 degrees) C) (3 squareroot 2, 135 degrees) D) (3 squareroot 2, 225 degrees)

OpenStudy (shamim):

the relation between polar coordinate n cartesian coordinate is\[r=\sqrt{x ^{2}+y ^{2}}\]

OpenStudy (anonymous):

thank you :) i just needed to know to do it

OpenStudy (shamim):

and\[\tan \theta =\frac{ y }{ x }\]

OpenStudy (anonymous):

there is only one choice there that matches... think of WHERE (-3, 3) is... what angle does that correspond to?

OpenStudy (anonymous):

wow i was overthinking it, thanks

OpenStudy (zzr0ck3r):

Is that right @jim_thompson5910

OpenStudy (zzr0ck3r):

nm

OpenStudy (zzr0ck3r):

its late

jimthompson5910 (jim_thompson5910):

r = sqrt(x^2 + y^2) r = sqrt((-3)^2 + (3)^2) r = sqrt((3)^2 + (3)^2) r = sqrt(2(3)^2) r = sqrt(3^2)*sqrt(2) r = 3*sqrt(2)

jimthompson5910 (jim_thompson5910):

theta = arctan(y/x) theta = arctan(3/(-3)) theta = arctan(-1) theta = -45 degrees theta = 315 degrees

OpenStudy (anonymous):

thank you guys i got it